# Role of defects in atom probe analysis of sol-gel silica

### v1

@@ 3.1 Experimental results
Figure 2: Mass spectra of silica samples, measured for the ranges (a) [0-80] and (b) [80-230] of the $m/z$ (Da) values with: Deep-UV (LEAP, dusty pink spectrum) and UV (LaWATAP, orange spectrum), 17 million ions collected at a laser energy of 250 pJ, corresponding to an energy density of 80 J/m2 for deep-UV, and at 50 nJ, corresponding to an energy density of 70 J/m2 for UV. The detection rate was set to 0.5% ions/pulse at $T=50$ K for deep-UV and to 0.25% ions/pulse at $T=50$ K for UV La-APT.
@@
The signal-to-noise ratio is better when a deep-UV light is used, with more Pt, Ga and C peaks being detected (see Figure 2b). We emphasise that the two analyses were not carried out under the same field conditions. To obtain an indication of the mean value of the electric field during the analyses, we calculated the charge state ratio (CSR) of silicon as the ratio of Si2+ ions to the total number of Si ions. According to Kingham’s theory, the CSR is an indicator of the electrostatic field in the entire analysis 42. It should be noted that the Kingham curves for silicon should be treated with caution for a material with properties that differ drastically from pure silicon and the values for the field should be regarded as estimates. In the deep-UV analysis, the mean value of the CSR was 0.6, which corresponds to a mean field of 19.7 V/nm, and in the UV analysis 0.95, which corresponds to a field of 21.5 V/nm. The evolution of the CSR (and thus the field) during the analyses is shown in Figure 6a,d for deep-UV and UV, respectively, and will be discussed later. We also note that the detection rates for the analyses in the deep-UV (0.5%) and UV (0.25%) range are comparable. When transitioning from deep-UV to UV light, the sample temperature must decrease to maintain a similar detection rate, while the electric field is increased from 19.7 to 21.5 V/nm. This is a first indication that the two illumination conditions heat the sample differently, with the heating being stronger when using the deep-UV light. Furthermore, we point out that the energy densities of deep-UV and UV light are similar (80 and 70 J/m2 respectively). From this we can conclude that the silica samples have a higher absorption coefficient for deep-UV light than for UV light. Therefore, when using deep-UV light, analyses can be performed at a static field that is lower than when using UV light, reducing the risk of sample breakage due to the mechanical stress caused by the static field.
We also notice that working with a lower field also reduces the background noise, as shown in Figure 2a,b for the deep-UV analysis, since the evaporation rate between laser pulses depends on the field when the same value for the base temperature (50 K) is set for both analyses. At these fixed temperature conditions, a reduction in the electric field essentially reduces the emission 43.
@@
In the reconstructed data shown in Figure 3a,b for deep-UV and UV light, respectively, clusters of carbon and platinum were consistently observed in the reconstructed samples. While this is to be expected during sample preparation, where a platinum-carbon mixture is used to protect the sample and attach it to the flat top posts with a GIS, it is usually observed at the beginning or end of the analysis. In this case, however, it is observed throughout the analysis, suggesting that the platinum mixture can fill the micropores in this type of material during sample preparation. It is hypothesised that the precursor gas may diffuse into the pores of the material when the lift-out is welded to the flat top posts. Sodium is also detected throughout the analysis, which is to be expected in the synthesis of silica. The observed distribution of sodium shows a slightly higher atomic concentration at the beginning of the analysis. However, we note that a similar redistribution of another alkali metal, namely lithium, is also observed in films of similar materials due to the electrostatic field 44 and such effects must be taken into account here, as sodium ions could also migrate through the porous material. In addition, the diffusion of Pt, Ga and C in the silica volume is more pronounced in deep-UV analysis, probably due to the stronger heating by the laser, which favours diffusion (or, if the silica area is small enough, the laser spot can directly illuminate and heat the Pt weld during APT analysis).
@@
To compare the mass resolution recorded in the UV and deep-UV APT analyses, we show in Figure 4a a zoom of the mass spectra from Figure 2a in the Si2+ peak region. The mass resolution is defined as the ratio between the width of the peak and the value of its maximum, where the width can be calculated at 10% and 1% of the maximum. The mass resolution at 10% is 240 for the deep-UV analysis and 160 for the UV analysis. However, at 1% the situation is reversed, with a better resolution for the UV APT. Indeed, the tails on the right side of the peaks in the UV spectrum are shorter. The peak tails have been shown to be related to the heating and cooling processes of the sample after laser illumination, causing a delay in the emission of the ions 12, 13. To analyse this process, it is better to look at the time-of-flight spectra in Figure 4b. Previous studies on La-APT have shown that the cooling time is a function of the heated area, the diffusivity of the sample and its geometry, in particular the cone angle 45, 12. The size of the heated area with deep-UV is less than 1 $\mu$m and almost the same size was measured and calculated for UV, taking into account the diffraction of the light at the apex of the sample 13. The different cooling times between the two analyses are therefore essentially due to the different geometries of the two samples. From the SEM images, we calculated an average cone angle of 6.5° and 15° for the deep-UV and UV samples, respectively. A change in the cone angle by almost a factor of 3 leads to a change in the cooling time by more than a factor of 10 due to heat propagation 45. The tails of the spectra in Figure 4b were fitted using the time-dependent evaporation rate equation 13:
$\Phi(t)=A\cdot\exp\Biggl{(}-\dfrac{Q}{k_{\mathrm{B}}\Biggl{(}T_{0}+\dfrac{\Delta T_{max}}{\sqrt{1+\frac{t-t_{0}}{\tau}}}\Biggr{)}}\Biggr{)},$ (2)
where $k_{\mathrm{B}}$ is the Boltzmann constant, $A=0.1\cdot{\exp\Biggl{(}{\frac{Q}{k_{B}(T_{0}+\Delta T_{max})}\Biggr{)}}}$ is a normalisation factor, $T_{0}$ is the initial temperature of the tip before laser illumination, which is fixed at 50 K, $\Delta T_{max}$ is the maximum temperature increase after laser illumination. The tip cooling process starts at $t=t_{0}$, defined as the point where the Si2+ peak intensity falls to 10% of its maximum, $\tau$ is the cooling time and $Q\approx Q_{0}(1-\frac{E}{E_{\mathrm{evap}}})$ is the height of the energy barrier, which depends linearly on the applied electric field $E$ as it approaches the evaporation field $E_{\mathrm{evap}}$. The values for the cooling time $\tau$ obtained by fitting the tail of the 28Si2+ peak in the time-of-flight spectrum in Figure 4b are 6 ns and 0.5 ns for the deep-UV and UV analyses 12. Further details on the fits can be found in the SI.
@@
Spikes in the detection rate were repeatedly observed in the mass spectra of both deep-UV and UV lasers, as shown in Figure 5a,b. These spikes can be attributed to micro-fractures in the sample, where a small piece breaks off due to the mechanical stress caused by the electric field, resulting in a burst of ion impacts to be detected. This phenomenon can also be partially explained by the presence of micropores in the sample. If a pore is present, the field to which the surface atoms surrounding the pore are exposed is higher because the effective radius of the sample is smaller. This can lead to a rapid increase in the evaporation rate of the ions. At the same time, the mass-to-charge ratio of Si2+ shifts to higher values, as shown in the mass spectrum history plot in Figure 5a,b. This behaviour is more pronounced in UV analysis, which is probably due to the voltage regulation used for deep-UV, which attenuates the fluctuations in the detection rate. The shift in the mass spectrum correlated with the evaporation rate has already been reported for materials with a large band gap and low electrical conductivity such as MgO, SiO2 and diamond 15, 16, 17, 46. This is due to the high resistivity of silica, which leads to a voltage drop across the sample that increases with increasing emission current. The voltage drop causes the ions to be emitted at a lower potential so that the mass-to-charge ratio of Si2+ in Figure 5a,b shifts to higher values.
@@
Figure 6: O/Si atomic ratio (black line) and charge state ratio (CSR, red line) as a function of depth of analysis in silica samples analysed with deep-UV (a) and UV (d) laser illumination. The evolution of the O/Si ratio (b) and CSR (c) as a function of laser energy under deep-UV illumination at 30 K (red dots) and 50 K (black dots). The evolution of the O/Si ratio (black dots, panels (e) and (f)) and the CSR (red dots, panels (e) and (f)) as a function of the depth of analysis in silica samples analysed with a UV laser, where the laser energy (blue line in panel (e)) or the DC voltage (blue line in panel (f)) was varied stepwise.
@@
The chemical composition of the oxides in the APT analysis has been shown to depend on the values of the electric field 47. To test this dependence, we calculated the charge state ratio (CSR) of silica (see red dots in Figure 6a,d). Recall that the changes in the CSR during the APT analysis give an indication of the relative changes in the electric field. When compared with the measured composition, it was found that the field in silica was lower in the deep-UV analysis and decreased when Ga, Pt and C were present. In addition, the change in the O/Si ratio exactly follows that of the field, which indicates that silica exhibits the same behaviour as other materials with a large band gap. In the UV analysis, the field remains almost constant, as does the composition with an O/Si ratio of about 1.4 (black dots in Figure 6d).
We have also investigated the dependence of the composition on the energy of the laser pulse. When using LEAP APT (deep-UV light), the analyses are performed at a constant detection rate, so that an increase in laser energy is always accompanied by a decrease in the applied voltage and thus the field. Figures 6b, c show the dependence of the O/Si ratio and the CSR on the laser energy. These two quantities decrease with increasing laser energy with almost the same gradient. This means that the dependence of the composition on the laser energy is negligible. For the UV analyses, it was possible to change the laser energy and the voltage separately. The results are shown in Figure 6e,f. When the laser energy was increased from 12 to 48 nJ, the O/Si ratio increased by 10%, but when the voltage (i.e. the CSR) was increased from 4 to 6 kV (CSR from 0.08 to 0.8), the O/Si ratio increased by a factor of 3 from 0.4 to 1.2, clearly showing the strong dependence of the composition on the field values. For both deep-UV and UV laser illumination, the measured composition deviates from the expected nominal composition (O/Si ratio of 2), but the measured ratio approaches this value at higher electric fields, which corresponds to a higher CSR of Si.
@@ 3.2 Theoretical results
The atomic distribution in Figure 3 shows that a number of impurities can be detected in both UV and deep-UV analyses, which can dramatically change the optical properties of silica. To obtain a convincing interpretation of the role of defects in the absorption properties of amorphous silica in the APT setup and to explain the different emission rates for the frequency ranges investigated in this work, we have determined optical absorption spectra by electronic structure calculations for three different DFT-optimised models of silica, representing relative minima in the potential energy landscape for a cubic box of 1.426 nm size with 192 atoms at a density of 66.2 atoms/nm3 (see Figure 7a), including a number of representative defects.
@@
Atomic evaporation in APT analysis takes place under static and external laser fields. The effect of strong static fields on the optical absorption properties of silica matrices was therefore initially investigated for a constant electric field of 20 V/nm applied in the horizontal direction. Note that the value of the field is close to the experimental values calculated from the CSR for the two analyses. We assume that the electric field only affects the electronic wave function, while the nuclei remain unaffected. The optical properties for each configuration were evaluated using DFT with the PBE exchange-correlation functional. We note that the 0 of the energy axis coincides with the top of the valence band in all calculated absorption spectra. A consistent red shift of the optical absorption spectrum was observed in all three samples compared to a zero electrostatic field (see Figure 7b), which leads to a slight reduction in the band gap of the amorphous silica. However, this shift is relatively small and did not significantly change the overall absorption properties. These results indicate that the applied static field does not fundamentally disrupt the optical response of amorphous silica, but rather causes subtle systematic perturbations.
@@
Figure 7: (a) SiO2 simulation supercells with $N=192$ atoms, corresponding to an edge of the box $L=1.42$ nm. There are no defects in these structures. The oxygen atoms are shown in red, the silicon atoms in yellow. (b) Influence of the static electric field: average of the imaginary part of the dielectric tensor along the orthogonal Cartesian directions for samples 1 (left in the panel (a) above), 2 (centre), 3 (right). Pristine amorphous silica matrix without electric field (violet, orange and green lines) and with a constant electric field of 20 V/nm along the $x$ direction (black, blue and red lines). (c) Amorphous silica matrix with a non-bridging oxygen defect without electric field (orange line) and with a constant electric field of 20 V/nm along the $x$ direction (violet line). These calculations were performed with a PBE functional that notoriously underestimates the fundamental gap.
@@
Motivated by this indication, we investigated a large number of typical defects in the silica matrix, which can also occur during the sol-gel and lift-out process, in order to interpret the effective laser light absorption in the UV and deep-UV range. In particular, we have focussed on interstitial C, H, O and Na atoms as well as interstitial and substitutional gallium atoms with or without, near or far oxygen vacancies. The optical properties for each configuration were evaluated using DFT with the HSE exchange-correlation functional, as it reproduces the experimental bandgap of amorphous silica much more accurately than PBE (see Figure S.6 in the SI).
@@
From the analysis of the absorption spectra in Figure 8, we conclude that substitutional carbon atoms in different positions do not significantly affect the absorption properties of the silica matrix. However, interstitial carbon defects at different concentrations and coordination numbers increase the intragap light absorption of silica between 4 and 8 eV, which could explain the stronger absorption observed experimentally in the deep-UV range.
Figure 9: Absorption spectra of the silica samples shown in panel (a), for (b) interstitial hydrogen bound to oxygen (H def.) and (c) for an oxygen vacancy (O def.) in different configurations compared to the defect-free amorphous structures (no def.).
@@
We have also analysed the effects of the likely presence of atomic and molecular sodium as well as gallium from the lift-out process in the silica matrix, even in the presence of oxygen vacancies. Previous atomistic computer simulations provided a comprehensive understanding of the structure of sodium silicate glasses and showed that sodium atoms tend to cluster in alkali-rich regions 48, 49, which we modelled by a sodium molecule within the matrix (see left side of Figure 10a), and to aggregate around non-bridging oxygen atoms, which we modelled by a sodium molecule inside the matrix near oxygen vacancies (see left side of Figure 10b).
@@
The presence of interstitial atomic or molecular sodium does not significantly change the absorption spectra of the silica matrices (see right side of Figure 10a). In fact, we could not observe any defect states within the optical gap for interstitial sodium atoms, as the absorption spectra are essentially the mere overlap of the spectra of atomic (blue or green lines in the right panel of Figure 10a) or molecular sodium (red line in the right panel of Figure 10a) and the silica matrix. By introducing bridging oxygen vacancies into the matrix (see left side of Figure 10b), however, the sodium cluster aggregates near these vacancies and we observed an increase in absorption at about 4-6 eV (see right side of Figure 10b), which is in the range of the experimental results with deep-UV light at about 4.8 eV.
@@
Figure 11: (a) Absorption spectra of different silica samples with different numbers of interstitial gallium atoms, which may be close to or far from the oxygen vacancy. (b) Absorption spectra of different silica samples with different numbers of interstitial gallium atoms without oxygen defects, at least in the vicinity. (c) Absorption spectra of silica samples with two substitutional gallium atoms (blue line) also in the presence of interstitial gallium atoms (green line). (d) Model of a silica sample with the simultaneous presence of several interstitial gallium atoms (green), one and two oxygen vacancies (red); (e) model of a silica sample with the simultaneous presence of two substitutional (blue) and one interstitial (green) gallium atom.
@@
Finally, we have analysed the change in light absorption properties due to the appearance of interstitial (see green atoms in Figure 11d) and substitutional (see blue atoms in Figure 11e) gallium atoms, also in the presence of bridging oxygen vacancies. Our simulations show that the substitution of silicon by gallium atoms in the silica matrix leads to the formation of metallic states at the Fermi level (see green and blue lines around 0 eV in Figure 11c), which makes the material conductive. This behaviour is in stark contrast to experimental observations in which gallium-doped silica retains its insulating character. Consequently, a simple substitution model for the incorporation of gallium appears to be an unlikely representation of the actual atomic configuration in our samples.

### v2

@@ 3.1 Experimental results
Figure 2: Mass spectra of silica samples, measured for the ranges (a) [0-80] and (b) [80-230] of the $m/z$ (Da) values with: Deep-UV (LEAP, dusty pink spectrum) and UV (LaWATAP, orange spectrum); 17 million ions were collected at a laser energy of 250 pJ, corresponding to an energy density of 80 J/m2 for deep-UV, and at 50 nJ, corresponding to an energy density of 70 J/m2 for UV. The detection rate was set to 0.5% ions/pulse at a base temperature of $T=50$ K for deep-UV and to 0.25% ions/pulse at a base temperature of $T=50$ K for UV La-APT.
@@
The signal-to-noise ratio is better when a deep-UV light is used, with more Pt, Ga and C peaks being detected (see Figure 2b). We emphasise that the two analyses were not carried out under the same field conditions. To obtain an indication of the mean value of the electric field during the analyses, we calculated the charge state ratio (CSR) of silicon as the ratio of Si2+ ions to the total number of Si ions. According to Kingham’s theory, the CSR is an indicator of the electrostatic field in the entire analysis 51. It should be noted that the Kingham curves for silicon should be treated with caution for a material with properties that differ drastically from pure silicon and the values for the field should be regarded as estimates. In the deep-UV analysis, the mean value of the CSR was 0.6, which corresponds to a mean field of 19.7 V/nm, and in the UV analysis 0.95, which corresponds to a field of 21.5 V/nm. The evolution of the CSR (and thus the field) during the analyses will be discussed later. We also note that the detection rates for the analyses in the deep-UV (0.5%) and UV (0.25%) range are comparable (recall that in APT analysis, the detection rate can vary by several orders of magnitude, from 0.001% to 100%. In addition, the detection rate varies exponentially with the analysis parameters such as the applied voltage and laser energy. Therefore, when setting an automatic detection rate, it is difficult to control its fluctuations with an accuracy of a factor of 2). When transitioning from deep-UV to UV light, the sample temperature must decrease to maintain a similar detection rate, while the electric field is increased from 19.7 to 21.5 V/nm. This is a first indication that the two illumination conditions heat the sample differently, with the heating being stronger when using the deep-UV light. Furthermore, we point out that the energy densities of deep-UV and UV light are similar (80 and 70 J/m2 respectively). From this we can conclude that the silica samples have a higher absorption coefficient for deep-UV light than for UV light. Therefore, when using deep-UV light, analyses can be performed at a static field that is lower than when using UV light, reducing the risk of sample breakage due to the mechanical stress caused by the static field.
We also notice that working with a lower field also reduces the background noise, as shown in Figure 2a,b for the deep-UV analysis, since the evaporation rate between laser pulses depends on the field when the same value for the base temperature (50 K) is set for both analyses. At these fixed temperature conditions, a reduction in the electric field essentially reduces the emission 52.
@@
In the reconstructed data shown in Figure 3a,b for deep-UV and UV light, respectively, clusters of carbon and platinum were consistently observed in the reconstructed samples. While this is to be expected during sample preparation, where a platinum-carbon mixture is used to protect the sample and attach it to the flat top posts with a GIS, it is usually observed at the beginning or end of the analysis. In this case, however, it is observed throughout the analysis, suggesting that the platinum mixture can fill the micropores in this type of material during sample preparation. It is hypothesised that the precursor gas may diffuse into the pores of the material when the lift-out is welded to the flat top posts. Sodium is also detected throughout the analysis, which is to be expected in the synthesis of silica. The observed distribution of sodium shows a slightly higher atomic concentration at the beginning of the analysis. However, we note that a similar redistribution of another alkali metal, namely lithium, is also observed in films of similar materials due to the electrostatic field 53 and such effects must be taken into account here, as sodium ions could also migrate through the porous material. In addition, the diffusion of Pt, Ga and C in the silica volume is more pronounced in deep-UV analysis, probably due to the stronger heating by the laser, which favours diffusion (or, if the silica area is small enough, the laser spot can directly illuminate and heat the Pt weld during APT analysis).
@@
To compare the mass resolving power recorded in the UV and deep-UV APT analyses, we show in Figure 4a a zoom of the mass spectra from Figure 2a in the Si2+ peak region. The mass resolving power is defined as the ratio between the width of the peak and the value of its maximum, where the width can be calculated at 10% and 1% of the maximum. The mass resolving power at 10% is 240 for the deep-UV analysis and 160 for the UV analysis. However, at 1% the situation is reversed, with a better resolution for the UV APT. Indeed, the tails on the right side of the peaks in the UV spectrum are shorter. The peak tails have been shown to be related to the heating and cooling processes of the sample after laser illumination, causing a delay in the emission of the ions 13, 14. To analyse this process, it is better to look at the time-of-flight spectra in Figure 4b. Previous studies on La-APT have shown that the cooling time is a function of the heated area, the diffusivity of the sample and its geometry, in particular the cone angle 54, 13. The size of the heated area with deep-UV is less than 1 $\mu$m and almost the same size was measured and calculated for UV, taking into account the diffraction of the light at the apex of the sample 14. The different cooling times between the two analyses are therefore essentially due to the different geometries of the two samples. From the SEM images, we calculated an average cone angle of 6.5° and 15° for the deep-UV and UV samples, respectively. A change in the cone angle by almost a factor of 3 leads to a change in the cooling time by more than a factor of 10 due to heat propagation 54. The tails of the spectra in Figure 4b were fitted using the time-dependent evaporation rate equation 14:
$\Phi(t)=A\cdot\exp\Biggl{(}-\dfrac{Q}{k_{\mathrm{B}}\Biggl{(}T_{0}+\dfrac{\Delta T_{\mathrm{max}}}{\sqrt{1+\frac{t-t_{0}}{\tau}}}\Biggr{)}}\Biggr{)},$ (2)
where $k_{\mathrm{B}}$ is the Boltzmann constant, $A=0.1\cdot{\exp\Biggl{(}{\frac{Q}{k_{\mathrm{B}}(T_{0}+\Delta T_{\mathrm{max}})}\Biggr{)}}}$ is a normalisation factor, $T_{0}$ is the base temperature of the tip before laser illumination, which is fixed at 50 K, $\Delta T_{\mathrm{max}}$ is the maximum temperature increase after laser illumination. The tip cooling process starts at $t=t_{0}$, defined as the point where the Si2+ peak intensity falls to 10% of its maximum, $\tau$ is the cooling time and $Q\approx Q_{0}(1-\frac{E}{E_{\mathrm{evap}}})$ is the height of the energy barrier, which depends linearly on the applied electric field $E$ as it approaches the evaporation field $E_{\mathrm{evap}}$. The values for the cooling time $\tau$ obtained by fitting the tail of the 28Si2+ peak in the time-of-flight spectrum in Figure 4b are 6 ns and 0.5 ns for the deep-UV and UV analyses 13. Further details on the fits can be found in the SI (see Figure S.6).
@@
Spikes in the detection rate were repeatedly observed in the mass spectra of both deep-UV and UV lasers, as shown in Figure 5a,b. These spikes can be attributed to micro-fractures in the sample, where a small piece breaks off due to the mechanical stress caused by the electric field, resulting in a burst of ion impacts to be detected. This phenomenon can also be partially explained by the presence of mesopores in the sample. If a pore is present, the field to which the surface atoms surrounding the pore are exposed is higher because the effective radius of the sample is smaller. This can lead to a rapid increase in the evaporation rate of the ions. At the same time, the mass-to-charge ratio of Si2+ shifts to higher values, as shown in the mass spectrum history plot in Figure 5a,b. This behaviour is more pronounced in UV analysis, which is probably due to the voltage regulation used for deep-UV, which attenuates the fluctuations in the detection rate. The shift in the mass spectrum correlated with the evaporation rate has already been reported for materials with a large band gap and low electrical conductivity such as MgO, SiO2 and diamond 16, 17, 18, 55. This is due to the high resistivity of silica, which leads to a voltage drop across the sample that increases with increasing emission current. The voltage drop causes the ions to be emitted at a lower potential so that the mass-to-charge ratio of Si2+ in Figure 5a,b shifts to higher values.
@@
Figure 6: O/Si atomic ratio (black line) and charge state ratio (CSR, red line) as a function of depth of analysis in silica samples analysed with deep-UV (a) and UV (d) laser illumination. The evolution of the O/Si ratio (b) and the CSR (c) as a function of laser energy under deep-UV illumination at base temperatures of 30 K (red dots) and 50 K (black dots). The evolution of the O/Si ratio (black dots, panels (e) and (f)) and the CSR (red dots, panels (e) and (f)) as a function of the depth of analysis in silica samples analysed with a UV laser, where the laser energy (blue line in panel (e)) or the DC voltage (blue line in panel (f)) was varied stepwise.
@@
The chemical composition of the oxides in the APT analysis has been shown to depend on the values of the electric field 56. To test this dependence, we calculated the charge state ratio (CSR) of silica (see red dots in Figure 6a,d). Recall that the changes in the CSR during the APT analysis give an indication of the relative changes in the electric field. When compared with the measured composition, it was found that the field in silica was lower in the deep-UV analysis and decreased when Ga, Pt and C were present. In addition, the change in the O/Si ratio exactly follows that of the field, which indicates that silica exhibits the same behaviour as other materials with a large band gap. In the UV analysis, the field remains almost constant, as does the composition with an O/Si ratio of about 1.4 (black dots in Figure 6d).
We have also investigated the dependence of the composition on the energy of the laser pulse. When using LEAP APT (deep-UV light), the analyses are performed at a constant detection rate, so that an increase in laser energy is always accompanied by a decrease in the applied voltage and thus the field. Figure 6b,c show the dependence of the O/Si ratio and the CSR on the laser energy. These two quantities decrease with increasing laser energy with almost the same gradient. This means that the dependence of the composition on the laser energy is negligible. For the UV analyses, it was possible to change the laser energy and the voltage separately. The results are shown in Figure 6e,f. When the laser energy was increased from 12 to 48 nJ, the O/Si ratio increased by 10%, but when the voltage (i.e. the CSR) was increased from 4 to 6 kV (CSR from 0.08 to 0.8), the O/Si ratio increased by a factor of 3 from 0.4 to 1.2, clearly showing the strong dependence of the composition on the field values. For both deep-UV and UV laser illumination, the measured composition deviates from the expected nominal composition (O/Si ratio of 2), but the measured ratio approaches this value at higher electric fields, which corresponds to a higher CSR of Si.
@@ 3.2 Theoretical results
The atomic distribution in Figure 3 shows that a number of impurities can be detected in both UV and deep-UV analyses, which can dramatically change the optical properties of silica 57, 58. In this context, we note that, in principle, three effects can interact to reduce the band gap: firstly, the presence of a strong electrostatic field. Secondly, phonon excitation and/or lattice distortion due to the heating of the sample during irradiation. The bandgap energy of semiconductors or insulators such as silica, decreases with increasing temperature as the amplitude of atomic vibrations increases, leading to larger interatomic distances. In this work we only perform static single-point calculations of the absorption spectra on optimised geometries with defects, i.e. the nuclei are clamped and the strong static and time-dependent fields acting on the samples, which in principle have an influence on the change of the equilibrium bond length due to the ionic charges, only affect the electronic degrees of freedom. Finally, the presence of defects, which typically lead to the generation of intragap states. In particular, we analyse two types of defects: Substitutional defects, in which an atom of the silica matrix, be it Si or O, is replaced by another type of atom, such as carbon or gallium, but occupies the same position; and interstitial defects, i.e. crystal defects in which an atom of (the same or) a different type occupies an interstitial site in the crystal structure (e.g. interstitial gallium defects are gallium atoms that occupy a site in the crystal structure of silica where normally no atom is present).
@@
To obtain a convincing interpretation of the role of defects in the absorption properties of amorphous silica in the APT setup and to explain the different emission rates for the frequency ranges investigated in this work, we have determined the ab initio optical absorption spectra for three different DFT-optimised models of pure amorphous silica (see Figure 7a), all characterised by a cubic box of 1.426 nm size with 192 atoms at a density of 66.2 atoms/nm3 and representing relative minima in the potential energy landscape. Figures S.7 and S.8 of the SI also show the real and imaginary parts of $\epsilon(\omega)$ and the refractive index, which lies between 1.5 and 1.7 in the energy range of interest, for three defect-free amorphous structures to which we add a number of representative defects each time.
@@
Atomic evaporation in APT analysis takes place under static and external laser fields. The effect of strong static fields on the optical absorption properties of silica matrices was therefore initially investigated for a constant electric field of 20 V/nm applied in the horizontal direction. Note that the value of the field is close to the experimental values calculated from the CSR for the two analyses. The optical properties for each configuration were evaluated using DFT with the PBE exchange-correlation functional. We note that the 0 of the energy axis coincides with the top of the valence band in all calculated absorption spectra. A consistent red shift of the optical absorption spectrum was observed in all three samples compared to a zero electrostatic field (see Figure 7b), which leads to a slight reduction in the band gap of the amorphous silica. However, this shift is relatively small and did not significantly change the overall absorption properties. These results indicate that the applied static field does not fundamentally disrupt the optical response of amorphous silica, but rather causes subtle systematic perturbations.
@@
Figure 7: (a) SiO2 simulation supercells with $N=192$ atoms, all corresponding to an edge of the box $L=1.42$ nm. There are no defects in these structures. The oxygen atoms are shown in red, the silicon atoms in yellow. (b) Influence of the static electric field: average of the imaginary part of the dielectric tensor along the orthogonal Cartesian directions for samples 1 (left in the panel (a) above), 2 (centre), 3 (right). Pristine amorphous silica matrix without electric field (violet, orange and green lines) and with a constant electric field of 20 V/nm along the $x$ direction (black, blue and red lines). (c) Amorphous silica matrix with a non-bridging oxygen defect without electric field (orange line) and with a constant electric field of 20 V/nm along the $x$ direction (violet line). These calculations were performed with a PBE functional that notoriously underestimates the fundamental gap.
@@
Motivated by this indication and driven by our experimental results, we investigated a variety of typical defects in the silica matrix, which can also occur during the sol-gel and lift-out process, to interpret the effective laser light absorption in the UV and deep-UV range. In particular, we have focussed on interstitial C, H, O and Na atoms as well as interstitial and substitutional gallium atoms with or without, near or far oxygen vacancies. The optical properties for each configuration were evaluated using DFT with the HSE06 exchange-correlation functional, as it reproduces the experimental bandgap of amorphous silica much more accurately than PBE (see Figure S.5 in the SI).
@@
From the analysis of the absorption spectra in Figure 8a,b we conclude that substitutional carbon atoms in different positions do not significantly affect the absorption properties of the silica matrix (see right panel in Figure 8a). This conclusion could also be drawn from the observation that the valence band edge of SiO2 consists mainly of non-bonding $O2p$ orbitals 48. Substitution with carbon does not drastically change the band gap, which is consistent with the calculated absorption spectra. However, interstitial carbon defects at different concentrations and coordination numbers increase the intragap light absorption of silica between 4 and 8 eV (see right panel in Figure 8b), which could explain the stronger absorption observed experimentally in the deep-UV range.
Figure 9: Absorption spectra of the silica samples shown in panel (a), for (b) interstitial hydrogen (H def., blue colour) bound to oxygen and (c) for an oxygen vacancy (O def., red colour) in different configurations compared to the defect-free amorphous structures of Figure 7 (no def.).
@@
We have also analysed the effects of the likely presence of atomic and molecular sodium as well as gallium from the lift-out process in the silica matrix, even in the presence of oxygen vacancies. Previous atomistic computer simulations provided a comprehensive understanding of the structure of sodium silicate glasses and showed that sodium atoms tend to cluster in alkali-rich regions 59, 60, which we modelled by a sodium molecule within the matrix (see left side of Figure 10a), and to aggregate around non-bridging oxygen atoms, which we modelled by a sodium molecule inside the matrix near oxygen vacancies (see left side of Figure 10b).
@@
The presence of interstitial atomic or molecular sodium does not significantly change the absorption spectra of the silica matrices (see right side of Figure 10a). In fact, we could not observe any defect states within the optical gap for interstitial sodium atoms, as the absorption spectra are essentially the mere overlap of the spectra of atomic (blue and green lines in the right panel of Figure 10a) or molecular sodium (red line in the right panel of Figure 10a) and the defect-free silica matrix (yellow line in the right panels of Figure 10a,b). However, by introducing bridging oxygen vacancies into the matrix (see left side of Figure 10b), the sodium cluster aggregates near these vacancies and we observed an increase in absorption at about 4-6 eV (see right side of Figure 10b), which is in the range of our experimental results with deep-UV light at about 4.8 eV.
@@
Figure 11: (a) Model of a silica sample with the simultaneous presence of several interstitial gallium atoms (green), one and two oxygen vacancies (red); (b) model of a silica sample with the simultaneous presence of two substitutional (blue) and one interstitial (green) gallium atom. (c) Absorption spectra of silica samples with two substitutional gallium atoms (blue line) also in the presence of interstitial gallium atoms (green line). (d) Absorption spectra of different silica samples with different numbers of interstitial gallium atoms without oxygen defects, at least in the vicinity. (e) Absorption spectra of different silica samples with different numbers of interstitial gallium atoms, which may be close to or far from the oxygen vacancy.
@@
Finally, we have analysed the change in light absorption properties due to the presence of interstitial (see green atoms in Figure 11a) and substitutional (see blue atoms in Figure 11b) gallium atoms, also in the presence of bridging oxygen vacancies. Our simulations show that the substitution of silicon by gallium atoms in the silica matrix leads to the formation of metallic states at the Fermi level (see green and blue lines around 0 eV in Figure 11c), which makes the material conductive. This behaviour is in stark contrast to experimental observations in which gallium-doped silica retains its insulating character. Consequently, a simple substitution model for the incorporation of gallium appears to be an unlikely representation of the actual atomic configuration in our samples.